OpenStudy (anonymous):

Susie wants to start making and selling bracelets. Her costs are the large sum of money that includes her initial investment and the cost of materials for each bracelet. She has already planned how much she is going to sell her bracelets for. Explain how Susie can set up a system of equations to predict when she will start to make a profit.

OpenStudy (anonymous):

@Ahsome

OpenStudy (danjs):

hello

OpenStudy (anonymous):

hi

OpenStudy (danjs):

Cost = Initial Investment + Materials

OpenStudy (danjs):

What is the profit defines as?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

do u have to make two quations

OpenStudy (danjs):

yes

OpenStudy (danjs):

We are looking for the point where "Susie starts to make a profit"

OpenStudy (danjs):

so from the above we want Profit = 0 = (selling revenue) - (materials cost)

OpenStudy (danjs):

When she starts to make a profit, at that instant, the materials cost and the amount sold will be the same, she will be at profit = 0, breaking even

OpenStudy (danjs):

get that?

OpenStudy (danjs):

So we need to define some variables....

OpenStudy (anonymous):

okay

OpenStudy (danjs):

So lets set up the first equation... Money generated = price per bracelet x number of bracelets sold

OpenStudy (danjs):

pick variables for each one

OpenStudy (anonymous):

m=pb*nb

OpenStudy (anonymous):

who is susie. is she sexy

OpenStudy (anonymous):

just kidding now lets start with this problem

OpenStudy (danjs):

ok, lets call each one by a single variable... how about this.. I(x) = income generated from selling x bracelets P = selling price of one bracelet X = # of bracelets sold Income generated = price per bracelet * bracelets sold I(x) = P*X

OpenStudy (danjs):

Now, we need an equation for the money going OUT, the costs Let C(x) = cost incurred from selling x bracelets

OpenStudy (anonymous):

okay

OpenStudy (danjs):

The cost to make bracelets = Materials per bracelet + initial investment

OpenStudy (danjs):

C(x) = cost to make x bracelets V = Initial Investment amount b = cost per bracelet x = number of bracelets sold C(x) = b*x + V

OpenStudy (danjs):

So those are the two equations, C(x) - cost to make x bracelets, and I(x) - Income made from selling x bracelets

OpenStudy (danjs):

THey are asking when she starts to make a profit, in other words, when is she at the Break Even point.

OpenStudy (anonymous):

this is starting to make sense :)

OpenStudy (anonymous):

really terrible at math

OpenStudy (danjs):

cost = income (breaking even) C(x) = I(x) \(\large\color{red}{ {\bf }(b*x) +V = (P*x) }\) I(x) = income generated from selling x bracelets P = selling price of one bracelet x = # of bracelets sold C(x) = cost to make x bracelets V = Initial Investment amount b = cost per bracelet x = number of bracelets sold

OpenStudy (danjs):

So with the variables we defined, She will start making money when the income coming in equals the cost going out. \(\large\color{blue}{ {\bf I(x) = C(x)} }\) Using both the equations with the variables we defined, you get the red equation above. With a little rearranging... \(\Huge\color{red}{ {\bf P*x~=~b*x~+~V} }\) For the number of bracelets sold, X, Susie will start to make a profit when the value of P*x equals the value of [b*x + V]. In other words, the money made from selling the bracelets at a certain price, will need to be equal to the cost of making that many bracelets plus her initial investment V.

OpenStudy (danjs):

I know it is a lot, let me know if you dont understand something.

OpenStudy (anonymous):

yes i do understand!!!! finally someone help me understand.

OpenStudy (anonymous):

okay i am trying but my computers acting up. but i will do so!!! You are a big help!!!!! thanks:)

OpenStudy (danjs):

no prob, let me know if ya need anything else whenever.

OpenStudy (anonymous):

you bet!!!!